HOW MANY RATIONAL NUMBERS ARE THERE BETWEEN 5 AND 14?
PLEASE I NEED HELP WITH THIS
THANK YOU SO MUCH I REALLY APPRECIATE ALL UR HELP! :)

Answer by mmm4444bot(95)

Hello There:
If we add 1/10 to 5, we get 5 + 1/10 which equals 51/10 (a rational number with a decimal representation of 5.1).
If we add 1/100, we get 501/100 which equals 5.01.
If we add 1/1000, we get 5001/1000 which equals 5.001.
We can continue this forever:
50001/10000 = 5.0001
500001/100000 = 5.00001
5000001/1000000 = 5.00001
Since this process can go on forever, and each time we are adding a smaller and smaller amount to 5, there are an infinite number of rational numbers between 5 and 14.
~ Mark